OpenStudy (anonymous):

use algebraic methods to determine the domain and range for the function from to defined by the given equation. a. x^2 – 6x –y =0 b. 4x^2 – xy +1 = 0 c. 5x^2(y) – x + 2 = 0

OpenStudy (anonymous):

use algebraic methods to determine the domain and range for the function from x to y defined by the given equation. a. x^2 – 6x –y =0 b. 4x^2 – xy +1 = 0 c. 5x^2(y) – x + 2 = 0

myininaya (myininaya):

(a) x^2-6x-y=0 we can solve this for x by using the quadratic formula \[x=\frac{6 \pm \sqrt{36-4(-y)}}{2}=\frac{6 \pm \sqrt{36+4y}}{2}\]

myininaya (myininaya):

what does y have to be in order for x to exist?

OpenStudy (anonymous):

I am not sure? this question is confusing

myininaya (myininaya):

can you have a negative under the square root?

OpenStudy (anonymous):

no

myininaya (myininaya):

you want 36+4y to be bigger than or equal to zero, correct?

OpenStudy (anonymous):

yes correct

myininaya (myininaya):

so solve this for y: 36+4y>=0 and you will have your range

OpenStudy (anonymous):

-9 >= y

OpenStudy (anonymous):

how do i write the range

myininaya (myininaya):

4y>=-36 so y>=-9

OpenStudy (anonymous):

what is the domain?

myininaya (myininaya):

solve for y

myininaya (myininaya):

to find range we solve for x to find domain we solve for y so solve the orignal equation for y

OpenStudy (anonymous):

so y >= -9 is the range or domain?

OpenStudy (anonymous):

are you still here?

OpenStudy (anonymous):

i think you are supposed to solve each of these for y and think of y as a function of x

OpenStudy (anonymous):

but there is something wrong here because you have these each set = 0

OpenStudy (anonymous):

that is what my project say

myininaya (myininaya):

y is range we got y>=-9

OpenStudy (anonymous):

ok fine. so first one is \[x^2-6x-y=0\] so \[y=x^2-6x\] as a function \[f(x)=x^2-6x\] domain all real numbers since it is a polynomial range is -9 to infinity

myininaya (myininaya):

no we find domain solve x^2-6x-y=0 for x and you will see the domain is all real numbers because x^2-6x is a polynomial

myininaya (myininaya):

now*

myininaya (myininaya):

not no lol

OpenStudy (anonymous):

LOL u guys are great help

OpenStudy (anonymous):

second one. \[4x^2-xy+1=0\] \[xy=4x^2+1\] \[y=\frac{4x^2+1}{x}\] domain all real numbers except 0

myininaya (myininaya):

all you have to do is solve both for x and y when you solve for x you can find the range when you solve for y you can find the domain

OpenStudy (anonymous):

so the second one does not have a range

OpenStudy (anonymous):

range is also all real numbers except 0 because the numerator is \[4x^2+1\] which cannot be 0

OpenStudy (anonymous):

no, range is just the possible y values. y can be anything except 0 since the numerator of your fraction is never 0

myininaya (myininaya):

range is [-4,4] you find once you solve for x

OpenStudy (anonymous):

oops really? i will try it.

myininaya (myininaya):

oops o did somehting wrong its (-inf,-4]U[4,inf)

OpenStudy (anonymous):

is this for the second problem

myininaya (myininaya):

the problem that we are now we are finding range

OpenStudy (anonymous):

how do you like that. you get \[y^2-16\]under the radical so \[y^2-16\geq 0\] who knew?

myininaya (myininaya):

lol satellitle

myininaya (myininaya):

you remember all that inverse crap right?

OpenStudy (anonymous):

oh of course. this thing is a rational function with a slant asymptote. so it does skip values, never between -4 and 4. must be late!

OpenStudy (anonymous):

so the range is (-inf, -4)U(4,inf)

OpenStudy (anonymous):

is the last one \[5x^2y-x+2=0\]?

OpenStudy (anonymous):

yes you have the range correct.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if i have the last one correct is is \[y=\frac{x-2}{5x^2}\] domain all numbers except 0

OpenStudy (anonymous):

and the range?

OpenStudy (anonymous):

honestly i am little confused by this. myininaya said solve for x. i can try it.

OpenStudy (anonymous):

oh ok!

OpenStudy (anonymous):

if i use the quadratic formula i see under the radical \[1-20y\] meaning \[1-20y\geq 0\] \[y\leq \frac{1}{20}\] would held if i had the picture of this thing to see if that makes sense.

OpenStudy (anonymous):

so i think i will stick to that answer, and go to bed.

myininaya (myininaya):

you want me to check your range? lol which one is this?

myininaya (myininaya):

ok the last one

OpenStudy (anonymous):

last one. i have confused myself.

myininaya (myininaya):

if i didnt make a mistake here is the range

OpenStudy (anonymous):

i must be tired if i cannot multiply 4 by 5 by 2! i best retire.

myininaya (myininaya):

yeah you were close then we also have y in the denominator and we know we cannot divide by 0 and when y is 0 the denominator is 0 so our range does not include y=0

OpenStudy (anonymous):

ok maybe but i am still skeptical. what do you get if x = 2?

myininaya (myininaya):

hmm you get (2,0)

OpenStudy (anonymous):

looks like if x = 2, then y = 0 original equation is \[5x^2y-x+2=0\] replace x by 2 get \[20y-2+2=0\] \[20y=0\] \[y=0\]

OpenStudy (anonymous):

i should switch back to windows and run maple, graph the thing and then i would be more convinced.

myininaya (myininaya):

well and also the quadratic formula just failes when a=0 in this case when 5yx^2=0 since y=0 a=0 so the quadratic formula won't work for that part of the answer

myininaya (myininaya):

so the range is (-inf,1/40]

OpenStudy (anonymous):

i am afraid i do not believe that either.

myininaya (myininaya):

i believe in it lol

OpenStudy (anonymous):

as x goes to infinity \[\frac{x-2}{5x^2}\] goes to zero

myininaya (myininaya):

right there is a horizontal asymptote

OpenStudy (anonymous):

let x = 1000000

OpenStudy (anonymous):

bet you get something closer to 0 than 1/40

OpenStudy (anonymous):

oh wait. maybe it is always negative. oops. ok possible

OpenStudy (anonymous):

ok i am beginning to believe it. may be always less than 1/40. tomorrow i graph !

myininaya (myininaya):

hey satelittle you know what would really convince you? i do

OpenStudy (anonymous):

maple

myininaya (myininaya):

you know how to find the maximum right? that derivative crap?

OpenStudy (anonymous):

ok here we go.

myininaya (myininaya):

so you do :) yay!

myininaya (myininaya):

f'(x)=[5x^2-10x(x-2)]/[(5x^2)]^2=[-5x^2+20x]/(25x^4) a critical numbers isfour f(4)=1/40

OpenStudy (anonymous):

well it is nice and late and i have had a few beers. but it looks like extreme value at 0 and 4

myininaya (myininaya):

well 0 doesn't exist in the orignal function

OpenStudy (anonymous):

0 is out, 4 gives.... holy carp, what do you know.

myininaya (myininaya):

lol

OpenStudy (anonymous):

learn something new every day. thanks.

myininaya (myininaya):

who needs maple

OpenStudy (anonymous):

me. use it all the time when i am not running linux

myininaya (myininaya):

well it is a nice friend to have

OpenStudy (anonymous):

and so to bed, before i make more of a fool of myself. take care

myininaya (myininaya):

later

myininaya (myininaya):

i'm going too gn